# Disciplined Biconvex Programming

### v1

@@ 3.4 Infeasible start
where $\nu>0$ is a penalty parameter. The problem (3.11) is always feasible, since by choosing sufficiently large $s$ and $t$, all constraints can be satisfied. Moreover, if the original biconvex problem (2.1) is feasible, then for sufficiently large $\nu$, applying ACS to (3.11) will yield a final point $(x^{\star},y^{\star},s^{\star},t^{\star})$, such that ${\bf 1{}}^{T}s^{\star}+{\lVert t^{\star}\rVert}_{1}=0$, i.e., $(x^{\star},y^{\star})$ is feasible and partially optimal for (2.1[NW06, SDU+17]. The full algorithm is given as follows.
@@ 4 Disciplined biconvex programming
The DBCP ruleset is heavily based on the DCP convex ruleset, which consists of a library of convex atomic functions, and a convex syntax ruleset that prescribes how these atomic functions may be composed to express (more complex) convex optimization problems [GBY06]. Specifically, all functions in a DCP-compliant problem must be formed as an expression consisting of variables, constants or parameters, and atomic functions. The sign, curvature, and monotonicity of each DCP expression can be determined recursively from those of its constituent parts, based on which the convexity of the overall problem can be verified by checking whether each function in the problem satisfies the DCP composition rules [GBY06, BV04]. We make an observation that is critical for extending DCP to DBCP: The basic arithmetic properties, in particular, the composition property, of biconvex functions, as discussed in §2.2, are compatible with the DCP ruleset. In other words, an optimization problem formed by verifiable biconvex expressions according to the DCP ruleset is still biconvex.
@@
To construct disciplined biconvex expressions, we first note that the product of expressions that are both nonconstant is prohibited in DCP, since the convexity of such expressions cannot be determined in general [GBY06]; however, most biconvex expressions that appear in practice are constructed through variable multiplications. Hence, we introduce the following product rule for DBCP:
@@
Note that the above DBCP product rules do not include constant/parameter-variable multiplications, since such expressions are already covered by the DCP product-free ruleset [GBY06]. The second rule is to prevent expressions like x * y, y * z, and z * x from appearing simultaneously in the same optimization problem, which would lead to cyclic interactions between the variables x, y, and z, where the biconvexity of the overall expression cannot be guaranteed. According to the definition and basic properties of the biconvex functions listed in §2.2, expressions formed according to the above DBCP product rule are guaranteed to be biconvex.
@@ 5 Implementation
In this section, we introduce our Python implementation of the DBCP framework, as an extension to CVXPY [DB16, AVDB18]. The corresponding open-source package, named dbcp, is available at
@@ 6.2 Bilinear logistic regression
We consider a bilinear logistic regression problem for binary classification [DCP07]. Suppose that we are given a dataset $(X_{i},y_{i})$, $i=1,\ldots,m$, where each sample consists of a feature matrix $X_{i}\in{\bf R{}}^{n\times k}$ and a binary label $y_{i}\in\{0,1\}$. Our goal is to construct a bilinear classifier $\hat{y}=1$ if $\mathop{\bf tr}(U^{T}XV)>0$, and $\hat{y}=0$ otherwise, where $U\in{\bf R{}}^{n\times r}$ and $V\in{\bf R{}}^{k\times r}$ are the bilinear logistic regression coefficients with a predefined (maximum) rank $r$, and $\mathop{\bf tr}(M)$ denotes the trace of some square matrix $M$. To fit a bilinear logistic regression model to the dataset, we would like to solve the following bilinear maximum likelihood estimation problem:
@@
In this example, we set $m=300$, $n=20$, $k=10$, and $r=5$. The dataset $(X_{i},y_{i})$, $i=1,\ldots,m$, is generated synthetically using the make_classification function from the scikit-learn library [PVG+11]. After calling prob.solve() with the proximal regularization weight lbd=1 and the termination threshold gap_tolerance=1e-4, we obtain a final point with an objective value $5\times 10^{-3}$.
@@ 6.3 $k$-means clustering
Suppose that we are given a set of data points $x_{i}\in{\bf R{}}^{n}$, $i=1,\ldots,m$, and we would like to cluster them into $k$ groups, using the $k$-means clustering method. This problem can be formulated as the following biconvex optimization problem [ZYHB25]:
@@ 6.4 Dictionary learning
We consider the sparse dictionary learning problem [AEB06], which aims to find a dictionary matrix $D\in{\bf R{}}^{m\times k}$ and a sparse code matrix $X\in{\bf R{}}^{k\times n}$, such that the data matrix $Y\in{\bf R{}}^{m\times n}$ can be well approximated by their product $DX$, while the matrix $X$ is sparse and the matrix $D$ has bounded Frobenius norm. The dictionary learning problem can be formulated as the following biconvex optimization problem:
@@ 6.5 Blind deconvolution
Blind deconvolution is a technique used to recover some sharp signal or image from a blurred observation when the blur itself is unknown [CVR14]. It jointly estimates both the original signal and the blur kernel, with some prior knowledge about their structures. Suppose that we are given a data vector $d\in{\bf R{}}^{m+n-1}$, which is the convolution of an unknown sparse signal $x\in{\bf R{}}^{n}$ and an unknown smooth vector $y\in{\bf R{}}^{m}$ with bounded $\ell_{\infty}$-norm (i.e., bounded largest entry). Additionally, we have the prior knowledge that both the vectors $x$ and $y$ are nonnegative. The corresponding blind deconvolution problem can be formulated as the following biconvex optimization problem:
@@
4obj = cp.sum_squares(conv(x, y) - d) + alpha_sp * cp.norm1(x) + alpha_sm * cp.sum_squares(cp.diff(y))
5constr = [cp.norm(y, "inf") <= beta]
6prob = BiconvexProblem(cp.Minimize(obj), [[x], [y]], constr)
@@
In this example, we consider the vectors $x\in{\bf R{}}^{120}$, $y\in{\bf R{}}^{40}$, and the $\ell_{\infty}$-norm bound $\beta=1$. The ground truth vectors were generated according to a similar example used by Shen et al. [SDU+17]. The ground truth and the recovered signals after calling prob.solve() with regularization parameters $\alpha_{\rm sp}=0.1$ and $\alpha_{\rm sm}=0.2$ are shown in figure 5.
@@ 6.6 Fitting input-output hidden Markov models
We consider the fitting problem of a logistic input-output hidden Markov model (IO-HMM) to some dataset [JAP24]. Suppose that we are given a dataset $(x(t),y(t))$, $t=1,\ldots,m$, where each sample consists of an input feature vector $x(t)\in{\bf R{}}^{n}$ and an output label $y(t)\in\{0,1\}$, generated from a $K$-state IO-HMM, according to the following procedure: Let $\hat{z}(t)\in\{1,\ldots,K\}$, $t=1,\ldots,m$, be the state label of the IO-HMM with initial state distribution $p_{\rm init}\in{\bf R{}}^{K}$ with ${\bf 1{}}^{T}p_{\rm init}=1$ and transition matrix $P_{\rm tr}\in{\bf R{}}^{K\times K}$ with $P_{\rm tr}{\bf 1{}}={\bf 1{}}$. At the time step $t$, the state label $\hat{z}(t)$ is sampled according to
@@
Given the dataset $(x(t),y(t))$, $t=1,\ldots,m$, we are interested in recovering the transition matrix $P_{\rm tr}$, the model parameters $\theta_{1},\ldots,\theta_{K}$, and the unobserved state labels $\hat{z}(1),\ldots,\hat{z}(m)$. Noticing that the transition matrix $P_{\rm tr}$ can be easily estimated from the state labels $\hat{z}(t)$, $t=1,\ldots,m$, we consider the following biconvex optimization problem for fitting the IO-HMM [ZYHB25]:

### v2

@@ 3.4 Infeasible start
where $\nu>0$ is a penalty parameter. The problem (3.11) is always feasible, since by choosing sufficiently large $s$ and $t$, all constraints can be satisfied. Moreover, if the original biconvex problem (2.1) is feasible, then for sufficiently large $\nu$, applying ACS to (3.11) will yield a final point $(x^{\star},y^{\star},s^{\star},t^{\star})$, such that ${\bf 1{}}^{T}s^{\star}+{\lVert t^{\star}\rVert}_{1}=0$, i.e., $(x^{\star},y^{\star})$ is feasible and partially optimal for (2.1[nocedal2006numerical, shen2017disciplined]. The full algorithm is given as follows.
@@ 4 Disciplined biconvex programming
The DBCP ruleset is heavily based on the DCP convex ruleset, which consists of a library of convex atomic functions, and a convex syntax ruleset that prescribes how these atomic functions may be composed to express (more complex) convex optimization problems [grant2006disciplined]. Specifically, all functions in a DCP-compliant problem must be formed as an expression consisting of variables, constants or parameters, and atomic functions. The sign, curvature, and monotonicity of each DCP expression can be determined recursively from those of its constituent parts, based on which the convexity of the overall problem can be verified by checking whether each function in the problem satisfies the DCP composition rules [grant2006disciplined, boyd2004convex]. We make an observation that is critical for extending DCP to DBCP: The basic arithmetic properties, in particular, the composition property, of biconvex functions, as discussed in §2.2, are compatible with the DCP ruleset. In other words, an optimization problem formed by verifiable biconvex expressions according to the DCP ruleset is still biconvex.
@@
To construct disciplined biconvex expressions, we first note that the product of expressions that are both nonconstant is prohibited in DCP, since the convexity of such expressions cannot be determined in general [grant2006disciplined]; however, most biconvex expressions that appear in practice are constructed through variable multiplications. Hence, we introduce the following product rule for DBCP:
@@
Note that the above DBCP product rules do not include constant/parameter-variable multiplications, since such expressions are already covered by the DCP product-free ruleset [grant2006disciplined]. The second rule is to prevent expressions like x * y, y * z, and z * x from appearing simultaneously in the same optimization problem, which would lead to cyclic interactions between the variables x, y, and z, where the biconvexity of the overall expression cannot be guaranteed. According to the definition and basic properties of the biconvex functions listed in §2.2, expressions formed according to the above DBCP product rule are guaranteed to be biconvex.
@@ 5 Implementation
In this section, we introduce our Python implementation of the DBCP framework, as an extension to CVXPY [diamond2016cvxpy, agrawal2018rewriting]. The corresponding open-source package, named dbcp, is available at
@@
Biconvex atoms.
Biconvex expressions in dbcp are created from fundamental atomic functions. Most of the atoms used in dbcp are inherited from CVXPY, such as inner product and elementwise multiplication. An example of the extended atoms that are not supported by CVXPY is the one-dimensional discrete convolution operation between two 1d-nonconstant expressions, which can be called by the user via dbcp.convolve. It has the same behavior as the convolve function in CVXPY, except that the first expression need not be a constant. The biconvex atom library is extensible, so that the other additional atoms can be added as necessary.
@@ 6.2 Bilinear logistic regression
We consider a bilinear logistic regression problem for binary classification [dyrholm2007bilinear]. Suppose that we are given a dataset $(X_{i},y_{i})$, $i=1,\ldots,m$, where each sample consists of a feature matrix $X_{i}\in{\bf R{}}^{n\times k}$ and a binary label $y_{i}\in\{0,1\}$. Our goal is to construct a bilinear classifier $\hat{y}=1$ if $\mathop{\bf tr}(U^{T}XV)>0$, and $\hat{y}=0$ otherwise, where $U\in{\bf R{}}^{n\times r}$ and $V\in{\bf R{}}^{k\times r}$ are the bilinear logistic regression coefficients with a predefined (maximum) rank $r$, and $\mathop{\bf tr}(M)$ denotes the trace of some square matrix $M$. To fit a bilinear logistic regression model to the dataset, we would like to solve the following bilinear maximum likelihood estimation problem:
@@
In this example, we set $m=300$, $n=20$, $k=10$, and $r=5$. The dataset $(X_{i},y_{i})$, $i=1,\ldots,m$, is generated synthetically using the make_classification function from the scikit-learn library [pedregosa2011scikit]. After calling prob.solve() with the proximal regularization weight lbd=1 and the termination threshold gap_tolerance=1e-4, we obtain a final point with an objective value $5\times 10^{-3}$.
@@ 6.3 $k$-means clustering
Suppose that we are given a set of data points $x_{i}\in{\bf R{}}^{n}$, $i=1,\ldots,m$, and we would like to cluster them into $k$ groups, using the $k$-means clustering method. This problem can be formulated as the following biconvex optimization problem [zhu2025multi]:
@@ 6.4 Dictionary learning
We consider the sparse dictionary learning problem [aharon2006k], which aims to find a dictionary matrix $D\in{\bf R{}}^{m\times k}$ and a sparse code matrix $X\in{\bf R{}}^{k\times n}$, such that the data matrix $Y\in{\bf R{}}^{m\times n}$ can be well approximated by their product $DX$, while the matrix $X$ is sparse and the matrix $D$ has bounded Frobenius norm. The dictionary learning problem can be formulated as the following biconvex optimization problem:
@@ 6.5 Blind deconvolution
Blind deconvolution is a technique used to recover some sharp signal or image from a blurred observation when the blur itself is unknown [chaudhuri2014blind]. It jointly estimates both the original signal and the blur kernel, with some prior knowledge about their structures. Suppose that we are given a data vector $d\in{\bf R{}}^{m+n-1}$, which is the convolution of an unknown sparse signal $x\in{\bf R{}}^{n}$ and an unknown smooth vector $y\in{\bf R{}}^{m}$ with bounded $\ell_{\infty}$-norm (i.e., bounded largest entry). Additionally, we have the prior knowledge that both the vectors $x$ and $y$ are nonnegative. The corresponding blind deconvolution problem can be formulated as the following biconvex optimization problem:
@@
4obj = cp.Minimize(
5 cp.sum_squares(convolve(x, y) - d)
6 + alpha_sp * cp.norm1(x)
7 + alpha_sm * cp.sum_squares(cp.diff(y)))
8constr = [cp.norm(y, "inf") <= beta]
9prob = BiconvexProblem(obj, [[x], [y]], constr)
@@
In this example, we consider the vectors $x\in{\bf R{}}^{120}$, $y\in{\bf R{}}^{40}$, and the $\ell_{\infty}$-norm bound $\beta=1$. The ground truth vectors were generated according to a similar example used by Shen et al. [shen2017disciplined]. The ground truth and the recovered signals after calling prob.solve() with regularization parameters $\alpha_{\rm sp}=0.1$ and $\alpha_{\rm sm}=0.2$ are shown in figure 5.
@@ 6.6 Fitting input-output hidden Markov models
We consider the fitting problem of a logistic input-output hidden Markov model (IO-HMM) to some dataset [jha2024active]. Suppose that we are given a dataset $(x(t),y(t))$, $t=1,\ldots,m$, where each sample consists of an input feature vector $x(t)\in{\bf R{}}^{n}$ and an output label $y(t)\in\{0,1\}$, generated from a $K$-state IO-HMM, according to the following procedure: Let $\hat{z}(t)\in\{1,\ldots,K\}$, $t=1,\ldots,m$, be the state label of the IO-HMM with initial state distribution $p_{\rm init}\in{\bf R{}}^{K}$ with ${\bf 1{}}^{T}p_{\rm init}=1$ and transition matrix $P_{\rm tr}\in{\bf R{}}^{K\times K}$ with $P_{\rm tr}{\bf 1{}}={\bf 1{}}$. At the time step $t$, the state label $\hat{z}(t)$ is sampled according to
@@
Given the dataset $(x(t),y(t))$, $t=1,\ldots,m$, we are interested in recovering the transition matrix $P_{\rm tr}$, the model parameters $\theta_{1},\ldots,\theta_{K}$, and the unobserved state labels $\hat{z}(1),\ldots,\hat{z}(m)$. Noticing that the transition matrix $P_{\rm tr}$ can be easily estimated from the state labels $\hat{z}(t)$, $t=1,\ldots,m$, we consider the following biconvex optimization problem for fitting the IO-HMM [zhu2025multi]: